# Transfer Learning in Robotics: An Upcoming Breakthrough? A Review of Promises and Challenges

### v2

@@ 4 Challenges and Promising Research Directions
The aforementioned examples highlight that knowledge can be transferred across several robots, tasks, and environments, thus highlighting the potential of transfer learning for robotics. However, several key questions falling under the areas what, how, and when to transfer remain to be answered to realize the full potential of transfer learning in robotics. In this section, we describe the key challenges that currently constitute roadblocks on the way to the future of transfer learning in robotics.
@@ 4.1 Abstraction Levels in Robotics
These cognitive skill levels can also be roughly identified in robotics, where they intrinsically correspond to different abstraction levels (see Figure 7). At the lowest learning level, a robot simply mimics the motion of a teacher without an explicit understanding of the underlying goal. If the teacher and the learner have similar embodiments, the task can simply be abstracted as a joint-level (positions, velocity, or acceleration) trajectory. However, in the case of different embodiments, transferring joint trajectories will result in very different end-effector trajectories. The task’s abstraction level can be increased by specifying, for example, end-effector trajectories and leveraging Cartesian trajectory controllers. To deal with changes in the environment, both the learning and abstraction levels need to be raised. For instance, transferring end-effector trajectories fails if obstacles are present in the environment. In this case, the task needs to be imitated instead of mimicked, i.e., the goal must be explicitly identified by the robot. The task can therefore be abstracted using, e.g., movement primitives (Ijspeert et al., 2002), thus allowing the specification of the key components of the imitated trajectory, e.g., the goal position, while leveraging robot skills such as collision avoidance, localization, and object detection to reproduce the task in different environments. In some cases, the physical capabilities of the teacher and the learner are very different, so the learner cannot achieve a demonstrated task by imitating the teacher. Instead, the learner must infer the goal from the demonstrated task and develop a strategy to achieve the same goal (Schaal, 1999). In other words, the transfer should be conducted at the higher abstraction level corresponding to achieving the goal specifications without imitating the teacher-specific actions. This corresponds to the emulation learning level. Finally, on an even higher abstraction level, the teacher should ideally give only high-level verbal instruction to the robot such as “open the drawer”, or “clean the room”. This requires the robot to have a skill set resembling that of agents with higher cognitive functions.
@@
It is important to notice that the abstraction level has a direct influence on the capabilities that are required for successful transfer across robots, environments, and tasks. In particular, transfer at a given abstraction level requires abilities ranging from the bottom of the robot capability stack onto the abilities of the current level (see Figure 7-right). For instance, transfer at the level of verbal instructions demands robots not only to have an abstract understanding of the word, but also to be endowed with task and motion planners, a set of robot skills, and low-level controllers to successfully execute the target task on the target robot in the target environment. In this context, recent advances in foundation models are a promising research direction to endow robots with emulated high-level cognitive capabilities (Bommasani et al., 2021; Ahn et al., 2022; Driess et al., 2023). Such foundation models generate semantic plans required to execute a target task based on language and on continuous information collected by the robot (e.g., images, state vectors). Driess et al. (Driess et al., 2023) proposed to address the correspondence problem between tasks at the highest level, i.e., from a semantic perspective, by combining a large language model with perceptual inputs in an embodied multimodal model. Transfer between robots, tasks, and environments is then achieved via a large amount of training data and by training the models on several robots, tasks, and environments simultaneously. In other words, the transfer comes — to some extent — “for free” thanks to the large scale of foundation models. Importantly, such transfer happens only at the highest level, i.e., at the level of semantic planning, while low-level policies and planners are assumed to be given. In other words, transfer is not tackled at lower levels. As a consequence, the difficulty of transfer, as well as the resulting performance, is highly dependent on the capability stack that is made available a priori for each robot. Moreover, training a (still limited) low-level capability stack from scratch, as done, e.g., in (Ahn et al., 2022), requires months of data collection and is not scalable in the long run.
@@ 4.2 Robotics Transformers
The subsequent RT-2 (Brohan et al., 2023a) is a vision-language-action model based on vision-language models (Chen et al., 2023; Driess et al., 2023) trained on web-scale data and tuned with robotic actions. The largest RT-2 consists of $55$ billions parameters. The increased performance of RT-2 compared to RT-1 and other adjusted baseline models (such as VC-1 (Majumdar et al., 2023), R3M (Nair et al., 2022), MOO (Stone et al., 2023)) is attributed to the vision-language backbone combining co-finetuning the pre-trained model jointly on robotics and web data, so that the model considers more abstract visual concepts as well as robot actions. Interestingly, the largest RT-2 model displays encouraging emergent capabilities, where the model is able to use the high-level concepts acquired from the web-scale data such as relative relations between objects to complete tasks that were not present in that form in the robotic dataset. However, these emerging capabilities only emerged in the largest models, which necessitate a complex cloud infrastructure to be deployed. Therefore, they are currently unsuitable for deployment on robotics platforms and self-sufficient autonomous systems. Moreover, the model is not able to produce motions that are not covered by the large robotics dataset. Furthermore, the size of the model ($55$ billions parameters) can slow the model inference down to $1$Hz.
@@ 4.3 Universal Representations
Moreover, the advent of large language models brought a new breed of representation models that have been rapidly applied in all areas of robotics, e.g., in planning (Shah et al., 2023; Huang et al., 2022), manipulation (Jiang et al., 2022; Ren et al., 2023; Khandelwal et al., 2022), and navigation (Lin et al., 2022; Parisi et al., 2022; Gadre et al., 2022). Such models are excellent candidates to harvest novel universal representations for robotics.
To be successfully leveraged in various robotic scenarios, universal representations should be expressive, while remaining simple enough to facilitate downstream applications. This is usually achieved via a dimensionality reduction process by extracting low-dimensional latent representations from data. While this latent space was usually assumed to be Euclidean, i.e., flat, recent works have shown the superiority of curved spaces — manifolds like hypersphere, hyperbolic spaces, symmetric spaces, and product of thereof — to learn representations of data exhibiting hierarchical or cyclic structures (Nickel and Kiela, 2017; Gu et al., 2019; López et al., 2021). For instance, the compositionality of visual scenes can be preserved via hyperbolic latent representations, thus improving downstream performance in point cloud analysis (Montanaro et al., 2022) and unsupervised visual representation learning (Ge et al., 2023). This suggests that rethinking inductive bias in the form of the geometry of universal representations may also be relevant for robotics applications and for transfer learning in robotics. For example, data associated with robotics taxonomies are better represented in hyperbolic spaces (Jaquier et al., 2022) and manipulation tasks encoded as graphs in the context of visual action planning (Lippi et al., 2023) may benefit from non-Euclidean representations.
@@ 4.6 Metrics for Transfer Learning in Robotics
Metrics to measure the transfer quality in robotics transfer learning settings remain overlooked in the community. However, such metrics are crucial to evaluate and compare the performance of transfer learning algorithms. Robotics adds an additional challenge to the problem of defining suitable metrics for transfer learning: Indeed, transfer learning in robotics can be seen as a three-part transfer problem consisting of transfer across robots, tasks, and environments.
@@
Several metrics have been defined and directly optimized to solve each of these sub-problems. In this context, domain adaptation received considerable attention from the machine learning community in recent years. When the distribution of the source and target domains can be reliably estimated, simple divergences, e.g., the Kullback-Leibler (KL) divergence (Kullback and Leibler, 1951) or the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) for labeled data, and the $H\Delta H$ divergence (Ben-David et al., 2010) for unlabeled data, provide a quantitative estimate of the domain transfer gap. In robotics, a large body of work focuses on the sim-to-real gap — or in other words, the reality gap — as a specific domain gap $G_{\mathcal{D}}$. The sim-to-real gap is often measured as the capacity of a realistic simulator to emulate the real world. Collins et al. (Collins et al., 2019b) quantified the reality gap by comparing simulated robot trajectories, e.g. using Pybullet (Coumans and Bai, 2016) or Mujoco (Todorov et al., 2012), with real-world trajectories captured by a motion capturing system. Importantly, the simulators accurately model kinematics, but generally struggle with dynamics of robots interacting with objects. Zhang et al. (Zhang et al., 2020) specifically focused on the sim-to-real gap in robotics and predicted the transfer performance of reinforcement learning policies using a probabilistic dynamics model. Limited attention was devoted to designing metrics for transfer learning across tasks or robots. In particular, the existing literature related to skill transfer learning in human-robot cooperation (Liu et al., 2020) does not agree on a specific skill transfer metric to measure the task transfer gap $G_{\mathcal{T}}$. Although various metrics have also been proposed in the context of motion retargeting, see, e.g., (Gielniak et al., 2013; Penco et al., 2018), quantifying the quality of retargeted motions, as well as robot transfer gap $G_{\mathcal{R}}$, generally remain open questions.
We contend that a suitable transferability metric $M$ for robotics should consider all three settings of transfer. For instance, this metric may be defined as a simple combination of individual metrics for robot, environment, and task transfer, e.g.,
$M=\lambda_{1}G_{\mathcal{R}}+\lambda_{2}G_{\mathcal{D}}+\lambda_{3}G_{\mathcal{T}},$
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are weights adjusting the individual metric influence. Such transfer metric has the potential to bootstrap the development of different transfer learning methods for robotics. In addition to transferability metrics, we believe that measures of the transfer gaps should be further investigated to decide when and what to transfer.
@@ 5 Conclusion: The Future of Transfer Learning in Robotics
The rise of transfer learning implies its potential to enable robots to leverage available knowledge to learn and master novel situations efficiently. In this paper, we aimed at unifying the concept of transfer learning in robotics via a novel taxonomy acting as a bedrock for future developments in the field. Building on the successes of transfer learning in robotics, we outlined relevant challenges that have to be solved to realize its full potential. It is important to highlight that these challenges intrinsically relate to the three fundamental questions of transfer learning, i.e., what, how, and when to transfer.
When to transfer refers to the conditions under which transfer should occur. Although this aspect has been overlooked, it is crucial to understand in which situations transferring knowledge is beneficial to avoid phenomena such as negative transfer. This is especially important in the context of robotics, where performance is directly linked to the efficiency of the robot performing tasks in the real world as well as the safety of the robot and its environment.
What to transfer refers to the knowledge or experience that should be transferred in a given setting. Answering this question is paramount for the success of transfer learning in robotics. We believe that linking the capabilities of our systems to different abstraction levels is key to make headway on this question. Moreover, metrics providing the quantification of the transfer gap and of the quality of transfer are crucial in this regard.
How to transfer relates to the methods, techniques, or strategies that are used to transfer a given form of knowledge or experience. Conceptually, we believe that a large potential lies in universal representations obtained e.g. via foundational models. Importantly, the obtained universal concepts must be grounded via sensory-motor experience in order to obtain truly intelligent embodied agents.

### v3

@@ 4 Challenges and Promising Research Directions
The aforementioned examples highlight that knowledge can be transferred across several robots, tasks, and environments, thus highlighting the potential of transfer learning for robotics. However, several key questions falling under the areas of identifying the similarities and differences across tasks, environments, and robot to single out what should be transferred and when remain to be answered to realize the full potential of transfer learning in robotics. In this section, we describe the key challenges that currently constitute roadblocks on the way to the future of transfer learning in robotics.
@@ 4.1 Abstraction Levels in Robotics
These cognitive skill levels can also be roughly identified in robotics, where they intrinsically correspond to different abstraction levels (see Figure 7). At the lowest learning level, a robot simply mimics the motion of a teacher without an explicit understanding of the underlying goal. If the teacher and the learner have similar embodiments, the task can simply be abstracted as a joint-level (positions, velocity, or acceleration) trajectory. However, in the case of different embodiments, transferring joint trajectories will result in very different end-effector trajectories. The task’s abstraction level can be increased by specifying, for example, end-effector trajectories and leveraging Cartesian trajectory controllers. To deal with changes in the environment, both the learning and abstraction levels need to be raised. For instance, transferring end-effector trajectories fails if obstacles are present in the environment. In this case, the task needs to be imitated instead of mimicked, i.e., the goal must be explicitly identified by the robot. The task can therefore be abstracted using, e.g., movement primitives (Ijspeert et al., 2002), thus allowing the specification of the key components of the imitated trajectory, e.g., the goal position, while leveraging robot skills such as collision avoidance, localization, and object detection to reproduce the task in different environments. In some cases, the physical capabilities of the teacher and the learner are very different, so the learner cannot achieve a demonstrated task by imitating the teacher. Instead, the learner must infer the goal from the demonstrated task and develop a strategy to achieve the same goal (Schaal, 1999). In other words, the transfer should be conducted at the higher abstraction level corresponding to achieving the goal specifications without imitating the teacher-specific actions. This corresponds to the emulation learning level. Finally, on an even higher abstraction level, the teacher should ideally give only high-level verbal instruction to the robot such as “open the drawer”, or “clean the room”. This requires the robot to have a skill set resembling that of agents with higher cognitive functions. Such capabilities have the potential to facilitate transfer in more challenging settings, such as dual- and triple-mode transfer.
@@
It is important to notice that the abstraction level has a direct influence on the capabilities that are required for successful transfer across robots, environments, and tasks. In particular, transfer at a given abstraction level requires abilities ranging from the bottom of the robot capability stack onto the abilities of the current level (see Figure 7-right). For instance, transfer at the level of verbal instructions demands robots not only to have an abstract understanding of the word, but also to be endowed with task and motion planners, a set of robot skills, and low-level controllers to successfully execute the target task on the target robot in the target environment. In this context, recent advances in foundational models are a promising research direction to endow robots with emulated high-level cognitive capabilities (Bommasani et al., 2021; Ahn et al., 2022; Driess et al., 2023). Such foundation models generate semantic plans required to execute a target task based on language and on continuous information collected by the robot (e.g., images, state vectors). Driess et al. (Driess et al., 2023) proposed to address the correspondence problem between tasks at the highest level, i.e., from a semantic perspective, by combining a large language model with perceptual inputs in an embodied multimodal model. Transfer between robots, tasks, and environments is then achieved via a large amount of training data and by training the models on several robots, tasks, and environments simultaneously. In other words, the transfer comes — to some extent — “for free” thanks to the large scale of foundation models. Importantly, such transfer happens only at the highest level, i.e., at the level of semantic planning, while low-level policies and planners are assumed to be given. In other words, transfer is not tackled at lower levels. As a consequence, the difficulty of transfer, as well as the resulting performance, is highly dependent on the capability stack that is made available a priori for each robot. Moreover, training a (still limited) low-level capability stack from scratch, as done, e.g., in (Ahn et al., 2022), requires months of data collection and is not scalable in the long run.
@@ 4.2 Robotics Transformers
The subsequent RT-2 (Brohan et al., 2023a) is a vision-language-action model based on vision-language models (Chen et al., 2023b; Driess et al., 2023) trained on web-scale data and tuned with robotic actions. The largest RT-2 consists of $55$ billions parameters. The increased performance of RT-2 compared to RT-1 and other adjusted baseline models (such as VC-1 (Majumdar et al., 2023), R3M (Nair et al., 2022), MOO (Stone et al., 2023)) is attributed to the vision-language backbone combining co-finetuning the pre-trained model jointly on robotics and web data, so that the model considers more abstract visual concepts as well as robot actions. Interestingly, the largest RT-2 model displays encouraging emergent capabilities, where the model is able to use the high-level concepts acquired from the web-scale data such as relative relations between objects to complete tasks that were not present in that form in the robotic dataset. However, these emerging capabilities only emerged in the largest models, which necessitate a complex cloud infrastructure to be deployed. Therefore, they are currently unsuitable for deployment on robotics platforms and self-sufficient autonomous systems. Moreover, the model is not able to produce motions that are not covered by the large robotics dataset. Furthermore, the size of the model ($55$ billions parameters) can slow the model inference down to $1$Hz.
@@ 4.3 Universal Representations
Moreover, the advent of large language models and visual language models brought a new breed of representation models that have been rapidly applied in all areas of robotics, e.g., in planning (Shah et al., 2023; Huang et al., 2022), manipulation (Jiang et al., 2022; Ren et al., 2023; Khandelwal et al., 2022), and navigation (Lin et al., 2022; Parisi et al., 2022; Gadre et al., 2022). Such models are excellent candidates to harvest novel universal representations for robotics.
Some large visual language models jointly account for different modalities by encoding them in a shared latent space. For instance, CLIP (Radford et al., 2021) is pre-trained on a large dataset of images and associated textual descriptions. The model maps both modalities into a shared latent space using a contrastive loss function. The idea of combining representations from different modalities into a shared latent space was also explored in robotics. For instance, Tatiya et al. (2020, 2023) learned a common latent space from haptic feature space of multiple robots. Knowledge from source robots was then transferred through the latent space to facilitate object recognition by a target robot. Lee et al. (2020b) considered specific encoders for RGB, depth, force-torque, and proprioception modalities, which were aggregated into a multimodal representation with a multimodal fusion model. This shared representation was shown to improve the sample efficiency of the manipulation policy for a peg insertion task. The case of missing modalities during inference time was considered in (Silva et al., 2020). A perceptual model of the world was trained by assuming that some modalities may not be available at all times. The approach can therefore compensate for missing or corrupted modalities during execution. Such joint representations are crucial to design universal representations for transfer.
To be successfully leveraged in various robotic scenarios, universal representations should be expressive, while remaining simple enough to facilitate downstream applications. This is usually achieved via a dimensionality reduction process by extracting low-dimensional latent representations from data. While this latent space was usually assumed to be Euclidean, i.e., flat, recent works have shown the superiority of curved spaces — manifolds like hypersphere, hyperbolic spaces, symmetric spaces, and product of thereof — to learn representations of data exhibiting hierarchical or cyclic structures (Nickel and Kiela, 2017; Gu et al., 2019; López et al., 2021). For instance, the compositionality of visual scenes can be preserved via hyperbolic latent representations, thus improving downstream performance in point cloud analysis (Montanaro et al., 2022) and unsupervised visual representation learning (Ge et al., 2023). This suggests that rethinking inductive bias in the form of the geometry of universal representations may also be relevant for robotics applications and for transfer learning in robotics. For example, data associated with robotics taxonomies are better represented in hyperbolic spaces (Jaquier et al., 2024) and manipulation tasks encoded as graphs in the context of visual action planning (Lippi et al., 2023) may benefit from non-Euclidean representations.
@@ 4.6 Metrics for Transfer Learning in Robotics
Two types of metrics are relevant for transfer learning in robotics, namely (i) metrics measuring the quality of the transfer, and (ii) metrics measuring the transfer gap. Metrics measuring the transfer quality aim at quantifying algorithmic performance and allows the comparison of the performances of different algorithms on the target space after transfer. In particular, they can also determine when transfer learning is useful. Metrics measuring the transfer gap measure the the discrepancy between the source and target spaces. Essentially, they provide a notion of how different the source and target robots, environments, and tasks are.
Transfer quality metrics are crucial to evaluate and compare the performance of transfer learning algorithms. As such, they are key to the development of novel transfer learning methods. Transfer quality in machine learning is commonly measured by comparing the performance achieved on the target task with or without transfer learning. In this case, the quality metric includes not only the final performance (asymptotic performance), but also the initial benefit of the transfer (jumpstart performance), the time to reach a predefined performance threshold (time to threshold), as well as the sensitivity to different hyperparameter settings (Taylor and Stone, 2009; Taylor et al., 2007). In addition to measuring the transfer quality, further analysis of the transfer process can be conducted by comparing the number of required sub-source tasks, the number of demonstrations (Barreto et al., 2018; Zhu et al., 2020b), or the required quality (e.g., suboptimal, expert, oracle) of the source space (Zhu et al., 2020a). Recently, Chen et al. (2023a) highlighted the importance of robust unsupervised evaluation metrics for domain adaptation. Such metrics should be independent of the training method, consistent across hyperparameters and models, and robust to adversarial attacks. Most of the aforementioned transfer quality metrics can and are in fact already used for transfer learning in robotics (Zhu et al., 2023).
@@
Transfer gap metrics provide a measure of the discrepency between the source and target spaces. Note that such metrics may also be used to measure performance in certain circumstances. Robotics adds an additional challenge to the problem of defining suitable transfer gap metrics for transfer learning: Indeed, transfer learning in robotics can be seen as a three-part transfer problem consisting of transfer across robots, tasks, and environments. Several metrics have been defined and directly optimized to solve each of these sub-problems. In this context, domain adaptation received considerable attention from the machine learning community in recent years. When the distribution of the source and target domains can be reliably estimated, simple divergences, e.g., the Kullback-Leibler (KL) divergence (Kullback and Leibler, 1951) the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) for labeled data, or the $H\Delta H$ divergence (Ben-David et al., 2010) MDD (Zhang et al., 2019), and SND (Saito et al., 2021) for unlabeled data, provide a quantitative estimate of the domain transfer gap. In robotics, a large body of work focuses on the sim-to-real gap — or in other words, the reality gap — as a specific domain gap $G_{\mathcal{D}}$. The sim-to-real gap is often measured as the capacity of a realistic simulator to emulate the real world. Collins et al. (Collins et al., 2019b) quantified the reality gap by comparing simulated robot trajectories, e.g. using Pybullet (Coumans and Bai, 2016) or Mujoco (Todorov et al., 2012), with real-world trajectories captured by a motion capturing system. Importantly, the simulators accurately model kinematics, but generally struggle with dynamics of robots interacting with objects. Zhang et al. (Zhang et al., 2020) specifically focused on the sim-to-real gap in robotics and predicted the transfer performance of reinforcement learning policies using a probabilistic dynamics model. Limited attention was devoted to designing transfer gap metrics for transfer learning across tasks or robots. In particular, the existing literature related to skill transfer learning in human-robot cooperation (Liu et al., 2020) does not agree on a specific skill transfer metric to measure the task transfer gap $G_{\mathcal{T}}$. Although various metrics have also been proposed in the context of motion retargeting, see, e.g., (Gielniak et al., 2013; Penco et al., 2018), quantifying the quality of retargeted motions, as well as robot transfer gap $G_{\mathcal{R}}$, generally remain open questions. We contend that a suitable transfer gap metric $G$ for robotics should consider all three settings of transfer. For instance, this metric may be defined as a simple combination of individual metrics for robot, environment, and task transfer, e.g.,
$G=\lambda_{1}G_{\mathcal{R}}+\lambda_{2}G_{\mathcal{D}}+\lambda_{3}G_{\mathcal{T}},$
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are weights adjusting the individual metric influence. Such transfer metric has the potential to bootstrap the development of different transfer learning methods for robotics, as it gives insights about the discrepancy of each mode (i.e., robot, environment, and tasks) from source to target space.
@@ 5 Conclusion: The Future of Transfer Learning in Robotics
The rise of transfer learning implies its potential to enable robots to leverage available knowledge to learn and master novel situations efficiently. In this paper, we aimed at unifying the concept of transfer learning in robotics via a novel taxonomy acting as a bedrock for future developments in the field. Building on the successes of transfer learning in robotics, we outlined relevant challenges that have to be solved to realize its full potential. It is important to highlight that these challenges intrinsically relate to determining what can and cannot be transferred. As illustrated in this paper, transfer learning in machine learning largely relies on identifying whether the distribution of data across two domains remain similar. Identifying similarities and differences across robotic tasks amounts to more than comparing distributions. We have emphasized the need to delineate similarities across tasks, environments, and robots in an effort to ease identification of commonalities and differences in each case. Automatically identify similarities across two situations in robotics relies importantly on spelling out how much prior knowledge on the physics of the robot and environment is provided. With advances in the use of foundational models, the amount of prior knowledge readily available may ease this transfer.